Your Christmas ski vacation was great, but it unfortunately ran a bit over budget. All is not lost: You just received an offer in the mail to transfer your $13,500 balance from your current credit card, which charges an annual rate of 21.3 percent, to a new credit card charging a rate of 11.9 percent.

How much faster could you pay the loan off by making your planned monthly payments of $300 with the new card? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Number of months

  

What if there was a 2 percent fee charged on any balances transferred? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Number of months
Answer #1

No. of months can be calculated using N function on a calculator

PV = -13,500, FV = 0, PMT = 300, I/Y = 21.3%/12 => Compute N = 91.12 months for existing card

For new card, I/Y = 11.9%/12 => Compute N = 59.90 months for new card

=> No. of months saved = 31.22 months

If there was 2 % fee, PV = -13,500 x (1 + 2%) = - 13,770 => Compute N = 61.54 months

No. of months saved = 29.58 months